OpenStudy (anonymous):

The price of a cup of coffee was 2.40 yesterday. Today, the price rose to 2.75 Find the percentage increase. Round your answer to nearest tenth

OpenStudy (anonymous):

First, we must find the fraction between the new price Cnew and the old price Cold: \[\frac{C_{new}}{C_{old}}=\frac{2.75}{2.40}=1.1458333...\] We can now convert it to a percentage by multiplying by 100: \[1.1458333...\times 100\%=114.58333...\%\] Next, the question asks for the increase, so we take the percentage that is an increase above 100% by subracting that 100% from our previous answer: \[114.58333...\%-100\%=14.58333...\%\] Finally, the question asks to round to the nearest tenth. We do that by seeing if the hundreds decimal place is above or below five: our hundreds decimal place is 8, so we round up: 14.58333 rounds up to 14.6. Therefore, our final answer is 14.6%